# Crossings and diffusion in Poisson driven marked random connection models

### v1

@@ 2.3 Transport via diffusions on singular geometric structures
Suppose that the triple $(\lambda,\nu,\varphi)$ satisfies Assumptions 2.3. We consider the RCM$(\rho,\nu,\varphi)$ where $\rho\geq\lambda$. We then consider the 1d diffusion (with unit diffusion coefficient) on the graph $\bar{\mathcal{C}}(\omega,\underline{u})$ given by the infinite cluster $\mathcal{C}(\omega,\underline{u})$ thought of as immersed in ${\mathbb{R}}^{d}$ with edges given by line segments (we take the segment $\overline{xy}$ from $x$ to $y$ instead of the pair $\{x,y\}$). We introduce the random measure $\mu(\omega,\underline{u}):=\sum_{\{x,y\}}\mathrm{d}s_{x,y}$, where the sum is over the edges of the infinite cluster and $\mathrm{d}s_{x,y}$ is the arc-length measure (i.e. the one-dimensional Hausdorff measure $\mathcal{H}^{1}$) on the segment $\overline{xy}$ from $x$ to $y$. We point out that in this section we use the term diffusion without referring to a particular stochastic process, but to the generalized differential operator $v\mapsto-{\rm div}\nabla_{\mu(\omega,\underline{u})}v$. We refer to [Zhikov06] for its definition (which anyway will not be used in the rest). One can study stochastic homogenization for this model with the approach of [Zhikov06]*Section 7, where the same problem is considered for the supercritical cluster of Bernoulli bond percolation.
@@ 3.1 Poisson–Boolean model
Under this assumption there exists $\lambda_{c}(\nu)\in(0,+\infty)$ with the following property: for $\rho<\lambda_{c}(\nu)$ the graph $G_{d}(\omega)$ has only bounded connected components $P_{\rho,\nu}$–a.s., while for $\rho>\lambda_{c}(\nu)$ the graph $G_{d}(\omega)$ has a unique infinite connected component $\mathcal{C}(\omega)$ $P_{\rho,\nu}$–a.s. (see [Gouere08], the existence of at most one infinite connected component follows also from [Chebunin24]*Theorem 12.1 since for any intensity, (2) is satisfied). By combining Theorem 2.4 with the results of [DembinNaN] we get the following.
@@
    There exist $c_{1},c_{2}>0$ such that, for all $\ell\geq 1$, it holds
@@
(i) We apply [DembinNaN]*Corollary 1.6. To this aim we take $n$ large to have $\nu([0,n])>0$. We write $\nu|_{[0,n]}$ for the the restriction of $\nu$ to $[0,n]$ and we write $\nu_{n}$ for the probability measure $\frac{1}{\nu([0,n])}\nu|_{[0,n]}$. Trivially the measure $\rho\mathcal{L}\otimes\nu|_{[0,n]}$ equals the measure $\nu([0,n])\rho\mathcal{L}\otimes\nu_{n}$. Hence, by [DembinNaN]*Corollary 1.6, we can fix $n$ large enough to have $\nu([0,n])\rho>\lambda_{c}(\nu_{n})$. Then, Theorem 2.4 gives a lower bound for the LR crossings of the Boolean model given by the RCM associated to the triple $(\nu([0,n])\rho,\nu_{n},\mathbbm{1}_{\mathcal{S}_{\rm B}})$, since the triple $(\lambda,\nu_{n},\mathbbm{1}_{\mathcal{S}_{\rm B}})$ satisfies Assumption 2.3 for any $\lambda\in(\lambda_{c}(\nu_{n}),\rho)$. On the other hand, since the measure $\rho\mathcal{L}\otimes\nu$ dominates the measure $\nu([0,n])\rho\mathcal{L}\otimes\nu_{n}$, it is simple to conclude by stochastic domination.
@@
(iii) The diffusion matrix $D(\rho)$ is well defined since the model fits to the general abstract setting of e.g. [Faggionato25]. In this sense, we stress that the additional condition $\int_{0}^{+\infty}\nu(dm)m^{d+2}<+\infty$ implies for both $k=0$ and $k=2$ that the expression $\sum_{x\in\hat{\omega}:x\sim 0}|x|^{k}$ has finite expectation w.r.t. the Palm distribution $P_{\rho,\nu}(\cdot|0\in\mathcal{C}(\omega))$, which is one of the assumptions of [Faggionato25] (the others can be easily checked). The proof that $D(\rho)$ is a strictly positive scalar matrix is obtained by combining Item (i) with Lemma 2.8 and [Faggionato25]*Corollary 2.7 (similarly to the proof of Theorem 2.10).
@@ 3.2 Mott variable range hopping
The random graph $G_{d}$ is included in the Poisson–Boolean model with deterministic radius $\zeta/2$ on the PPP with intensity $\rho$, so for $\rho>0$ small enough, $P_{\rho,\nu}$-a.s. $G_{d}$ has no infinite cluster. On the other hand, fixing some $x\in(-\zeta/2,\zeta/2)\cap\operatorname{supp}\nu$, we have that $G_{d}$ contains the Poisson–Boolean model with deterministic radius $(\zeta/2-|x|)/4$ on the points with marks in $[x-(\zeta/2-|x|)/4,x+(\zeta/2-|x|)/4]$ (which form a PPP with intensity $\rho\nu([x-(\zeta/2-|x|)/4,x+(\zeta/2-|x|)/4])>0$). Hence for $\rho$ large enough, $P_{\rho,\nu}$-a.s. $G_{d}$ has at least one infinite cluster. The conclusion then follows by stochastic domination. ∎
@@
If $\operatorname{supp}\nu$ intersects both $(-\zeta/2,0)$ and $[0,\zeta/2)$ but $\nu$ is not good, then we can write the PPP with law $P_{\rho,\nu}$ as superposition of independent PPPs with laws $P_{\rho_{-},\nu_{-}}$ and $P_{\rho_{+},\nu_{+}}$ where $\rho_{-}:=\rho\nu((-\infty,A_{-}])$, $\nu_{-}=\nu(\cdot|(-\infty,A_{-}])$, $\rho_{+}:=\rho\nu([A_{+},+\infty))$, $\nu_{+}=\nu(\cdot|[A_{+},+\infty))$. A.s. there is no edge in the RCM$(\rho,\nu,\varphi)$ between these two PPPs (and in particular $\lambda_{c}(\nu)=\min\{\lambda_{c}(\nu_{-}),\lambda_{c}(\nu_{+})\}$) and for each one, Assumption 2.3(iv) is verified by Lemma 3.4. In particular, one can apply our general approach to $P_{\rho_{-},\nu_{-}}$ and $P_{\rho_{+},\nu_{+}}$ separately.
@@
Theorem 3.7 is an immediate application of Theorems 2.4, 2.10 and 2.12 with $\varphi:=\mathbbm{1}_{\mathcal{S}_{\rm M}}$ and $\lambda\in(\lambda_{c},\rho)$, hence we omit the proof.
@@
Suppose that $\nu\big{(}(-\zeta/2,\zeta/2)\big{)}>0$. Then for any $\rho>\lambda_{c}(\nu)$ there exist $c_{1},c_{2}>0$ such that, for all $\ell\geq 1$, it holds
@@ 4.1 RCM driven by a PPP on $\mathbb{R}^{d}\times\mathbb{M}$
[…]
Let $G=(V,E)$ be a graph as above (i.e. $G=G(\omega,\underline{u})$), where $V$ and $E$ are respectively the vertex set and the edge set. If $\{t,t^{\prime}\}$ is an edge, we write $t\sim t^{\prime}$ and say that $t,t^{\prime}$ are adjacent. A path in $G$ is a string $(t_{1},t_{2},\dots,t_{n})$ such that $t_{k}\in V$ for all $k=1,2,\dots,n$ and $\{t_{k},t_{k+1}\}\in E$ for $k=1,2,\dots,n-1$. Given $A,B,D\subset{\mathbb{R}}^{d}\times{\mathbb{M}}$ we say that there is a path from $A$ to $B$ in $D$ if there is a path $(t_{1},t_{2},\dots,t_{n})$ in $G$ such that $t_{1}\in A$, $t_{n}\in B$ and $t_{k}\in D$ for all $k=1,2,\dots,n$. In this case we write “$A\leftrightarrow B$ in $D$” or also “$A\leftrightarrow B$ in $D$ in $G$” when we need to stress that we deal with the graph $G$. If $D={\mathbb{R}}^{d}\times{\mathbb{M}}$, it is omitted from the notation.
@@ 4.2 FKG inequality
To benefit of the FKG inequality we need to switch to a different space. As in [Last17], we define $({\mathbb{R}}^{d}\times{\mathbb{M}})^{[2]}$ as the set $\{A\subset{\mathbb{R}}^{d}\times{\mathbb{M}}\,:\,\sharp A=2\}$. When equipped with the Hausdorff metric this set is a Borel subset of a complete separable metric space [Last17]. We set ${\mathbb{W}}:=({\mathbb{R}}^{d}\times{\mathbb{M}})^{[2]}\times[0,1]$ (trivially, an element of ${\mathbb{W}}$ is a pair ($\big{(}\{t,t^{\prime}\},u\big{)}$ with $t\not=t^{\prime}$ in ${\mathbb{R}}^{d}\times{\mathbb{M}}$ and $u\in[0,1]$). We then define $\mathcal{N}({\mathbb{W}})$ as the family of locally finite subsets of ${\mathbb{W}}$. Then $\mathcal{N}({\mathbb{W}})$ is a measurable space with $\sigma$–algebra of measurable sets generated by the sets $\{\zeta\in\mathcal{N}({\mathbb{W}})\,:\,\sharp(\zeta\cap B)=n\}$ as $B$ varies among the Borel sets of ${\mathbb{W}}$ and $n$ varies in ${\mathbb{N}}$.
On ${\mathbb{W}}$, we define the partial order $\leq$ as follows: $\zeta\leq\zeta^{\prime}$ if for any ($(\{t,t^{\prime}\},u)\in\zeta$ it holds ($(\{t,t^{\prime}\},u^{\prime})\in\zeta^{\prime}$ for some $u^{\prime}\in[0,u]$. A measurable function $f:\mathcal{N}({\mathbb{W}})\to{\mathbb{R}}$ is called increasing if $f(\zeta)\leq f(\zeta^{\prime})$ whenever $\zeta\leq\zeta^{\prime}$. An event $A$ in $\mathcal{N}({\mathbb{W}})$ is called increasing if $\mathbbm{1}_{A}$ is increasing.
@@
Indeed, the first inequality follows from the FKG inequality for independent uniform random variables with the partial order naturally induced by the standard order on real numbers ($f$ and $g$ are both decreasing functions of the connection variables $u$). The second inequality follows from the FKG inequality for PPPs (cf. [Last18]*Theorem 20.4) as $E[f|\eta]$ and $E[g|\eta]$ are both increasing functions of $\eta$ w.r.t. inclusion (see below), while under $P$ $\eta$ is a PPP with law $P_{\rho,\nu}$. To check the last monotonicity, we observe that, given $\eta,\eta^{\prime}\in\Omega$ with $\eta\subset\eta^{\prime}$, we can realize a coupling between the connection variables associated to $\eta$ and the ones associated to $\eta^{\prime}$ so that for all $x\not=y$ in $\eta\subset\eta^{\prime}$ the connection variable for the pair $\{x,y\}$ is the same for $\eta$ and for $\eta^{\prime}$. Then, the corresponding configurations in $\mathcal{N}({\mathbb{W}})$ are in increasing order w.r.t. $\leq$ (where each configuration is given by the collections of triples ($(\{x,y\},u)$, $u$ being the connection variable associated to $\{x,y\}$). The same order is then maintained when applying $f$, or $g$. ∎
@@ 4.3 Irreducibility
For $P_{\lambda,\nu}\otimes{\mathbb{Q}}$–a.a. $(\omega,\underline{u})$ the following holds: for each $(x,m)\in\omega$ with $m\in A$ the cluster of $(x,m)$ in $G(\omega,\underline{u})$ is finite.
@@
By Lemma 4.2, given a marked PPP ${\mathbb{X}}$ with law $P_{\lambda,\nu}$, if we remove from $G({\mathbb{X}})$ all vertices $(x,m_{x})$ with $m_{x}\in A$ and the associated edges, the resulting graph has the same infinite clusters as $G({\mathbb{X}})$. By Assumption 2.3(v), we conclude that $\nu(A)<1$ with $A$ from (14). Set $\lambda^{\prime}=\lambda\nu(A^{c})$ and $\nu^{\prime}:=\nu(\cdot|A^{c})$. Let us verify that the triple $(\lambda^{\prime},\nu^{\prime},\varphi)$ satisfies Assumption 2.3 (iv) and (v). The graph obtained above by removing vertices $(x,m_{x})$ with $m_{x}\in A$ and projected on ${\mathbb{R}}^{d}$ is just the RCM$(\lambda^{\prime},\nu^{\prime},\varphi)$ and, by Assumption 2.3(v), it has an unbounded connected component $P_{\lambda^{\prime},\nu^{\prime}}\otimes Q$–a.s. Hence, the triple $(\lambda^{\prime},\nu^{\prime},\varphi)$ satisfies Assumption 2.3 (v).
@@
${\mathbb{P}}(t\leftrightarrow t^{\prime}\leftrightarrow\infty\text{ in }G({\mathbb{X}}\cup\{t,t^{\prime}\}))>0\qquad\forall(t,t^{\prime})\in({\mathbb{R}}^{d}\times A^{c})^{2}\,.$ (15)
@@
$f(t,t^{\prime},\eta):=\mathbbm{1}(t,t^{\prime}\in{\mathbb{R}}^{d}\times A^{c}){\mathbb{P}}\left({\mathbb{R}}^{d}\times A\leftrightarrow t\leftrightarrow t^{\prime}\leftrightarrow\infty\text{ in }G(\eta)\right)\,.$
@@
(we stress that now the expectation is w.r.t. both the PPP ${\mathbb{X}}$ and the connection variables). By the multivariate Mecke equation (cf. [Last18]*Theorem 4.4), we conclude that for $(\lambda\mathcal{L}\otimes\nu)^{\otimes 2}$ a.a. $(t,t^{\prime})\in({\mathbb{R}}^{d}\times A^{c})^{2}$ it holds ${\mathbb{E}}[f(t,t^{\prime},{\mathbb{X}}\cup\{t,t^{\prime}\})]=0$, which can be rewritten as ${\mathbb{P}}\left({\mathbb{R}}^{d}\times A\leftrightarrow t\leftrightarrow t^{\prime}\leftrightarrow\infty\text{ in }G({\mathbb{X}}\cup\{t,t^{\prime}\})\right)=0$. Combining this result with (15), we get ${\mathbb{P}}(t\leftrightarrow t^{\prime}\text{ in }G(({\mathbb{X}}\setminus({\mathbb{R}}^{d}\times A))\cup\{t,t^{\prime}\}))>0$ for $(\lambda\mathcal{L}\otimes\nu)^{\otimes 2}$–a.a. $(t,t^{\prime})\in({\mathbb{R}}^{d}\times A^{c})^{2}$. The last property is equivalent to (3) for $\lambda^{\prime},\nu^{\prime}$.
We have proved that the triple $(\lambda^{\prime},\nu^{\prime},\varphi)$ satisfies Assumption 2.3. Clearly, (13) is valid with $\nu$ and ${\mathbb{X}}$ replaced by $\nu^{\prime}$ and a marked PPP with law $P_{\lambda^{\prime},\nu^{\prime}}$, respectively. Since RCM$(\lambda^{\prime},\nu^{\prime},\varphi)$ can be thought of as a subgraph of RCM$(\lambda,\nu,\varphi)$, this completes the reduction of Theorem 2.4, showing that we can assume (13) w.l.o.g.

### v2

@@ 2.3 Transport via diffusions on singular geometric structures
[…]
@@ 3.1 Poisson–Boolean model
Under this assumption, there exists $\lambda_{c}(\nu)\in(0,+\infty)$ with the following property: for $\rho<\lambda_{c}(\nu)$ the graph $G_{d}(\omega)$ has only bounded connected components $P_{\rho,\nu}$–a.s., while for $\rho>\lambda_{c}(\nu)$ the graph $G_{d}(\omega)$ has a unique infinite connected component $\mathcal{C}(\omega)$ $P_{\rho,\nu}$–a.s. (see [Gouere08], the existence of at most one infinite connected component follows also from [Chebunin24]*Theorem 12.1 since for any intensity, (2) is satisfied). By combining Theorem 2.4 with the results of [DembinToappear] we get the following.
@@
    There exist $c_{1},c_{2}>0$ such that, for all $\ell$ large enough, it holds
@@
(i) We apply [DembinToappear]*Corollary 1.6. To this aim we take $n$ large to have $\nu([0,n])>0$. We write $\nu|_{[0,n]}$ for the the restriction of $\nu$ to $[0,n]$ and we write $\nu_{n}$ for the probability measure $\frac{1}{\nu([0,n])}\nu|_{[0,n]}$. Trivially the measure $\rho\mathcal{L}\otimes\nu|_{[0,n]}$ equals the measure $\nu([0,n])\rho\mathcal{L}\otimes\nu_{n}$. Hence, by [DembinToappear]*Corollary 1.6, we can fix $n$ large enough to have $\nu([0,n])\rho>\lambda_{c}(\nu_{n})$. Then, Theorem 2.4 gives a lower bound for the LR crossings of the Boolean model given by the RCM associated to the triple $(\nu([0,n])\rho,\nu_{n},\mathbbm{1}_{\mathcal{S}_{\rm B}})$, since the triple $(\lambda,\nu_{n},\mathbbm{1}_{\mathcal{S}_{\rm B}})$ satisfies Assumption 2.3 for any $\lambda\in(\lambda_{c}(\nu_{n}),\rho)$. On the other hand, since the measure $\rho\mathcal{L}\otimes\nu$ dominates the measure $\nu([0,n])\rho\mathcal{L}\otimes\nu_{n}$, it is simple to conclude by stochastic domination.
@@
(iii) The diffusion matrix $D(\rho)$ is well defined since the model fits to the general abstract setting of e.g. [Faggionato25]. In this sense, we stress that the additional condition $\int_{0}^{+\infty}\nu(dm)m^{d+2}<+\infty$ implies for both $k=0$ and $k=2$ that the expression $\sum_{x\in\hat{\omega}:x\sim 0}|x|^{k}$ has finite expectation w.r.t. the Palm distribution $P_{\rho,\nu}(\cdot|0\in\mathcal{C}(\omega))$, which is one of the assumptions of [Faggionato25] (the others can be easily checked). The proof that $D(\rho)$ is a strictly positive scalar matrix is obtained by combining Item (i) with Lemma 2.9 and [Faggionato25]*Corollary 2.7 (similarly to the proof of Theorem 2.11). ∎
@@ 3.2 Mott variable range hopping
The random graph $G_{d}$ is included in the Poisson–Boolean model with deterministic radius $\zeta/2$ on the PPP with intensity $\rho$, so for $\rho>0$ small enough, $P_{\rho,\nu}$-a.s. $G_{d}$ has no infinite cluster. On the other hand, fixing some $x\in(-\zeta/2,\zeta/2)\cap\operatorname{supp}\nu$, we have that $G_{d}$ contains the Poisson–Boolean model with deterministic radius $(\zeta/2-|x|)/4$ on the points with marks in $x-(\zeta/2-|x|)/4,x+(\zeta/2-|x|)/4]$ (which form a PPP with intensity $\rho\nu([x-(\zeta/2-|x|)/4,x+(\zeta/2-|x|)/4])>0$). Hence for $\rho$ large enough, $P_{\rho,\nu}$-a.s. $G_{d}$ has at least one infinite cluster. The conclusion then follows by stochastic domination. ∎
@@
[…]
@@
Theorem 3.7 is an immediate application of Theorems 2.4, 2.11 and 2.13 with $\varphi:=\mathbbm{1}_{\mathcal{S}_{\rm M}}$ and $\lambda\in(\lambda_{c},\rho)$, hence we omit the proof.
@@
Suppose that $\nu\big((-\zeta/2,\zeta/2)\big)>0$. Then for any $\rho>\lambda_{c}(\nu)$ there exist $c_{1},c_{2}>0$ such that, for all $\ell$ large enough, it holds
@@ 4.1 RCM driven by a PPP on $\mathbb{R}^{d}\times\mathbb{M}$
Given a configuration $\omega\in\Omega$, we can enumerate the points in $\omega$ in a measurable way [Last18]*Proposition 6.2, i.e. the map $\Omega\to({\mathbb{R}}^{d}\times{\mathbb{M}})\cup\{\partial\}:\omega\mapsto t_{i}(\omega)$ is measurable, where $t_{i}(\omega)$ is the ii–th point of $\omega$ if it exists, or the abstract point $\partial$ if the ii–th point does not exist. When we write $\omega=\{t_{1},t_{2},\dots\}$ we mean that $t_{1},t_{2},\dots$ is the resulting enumeration. Recall that $J:=\{(i,j)\in{\mathbb{N}}_{+}\times{\mathbb{N}}_{+}\,:\,1\leq i<j\}$. Given $\omega=\{t_{1},t_{2},\dots\}\in\Omega$ and $\underline{u}\in[0,1]^{J}$, the graph $G(\omega,\underline{u})$ is defined as follows: it has vertex set $\omega$ and edge set given by the pairs $\{t_{i},t_{j}\}$ of distinct vertices such that $u_{i,j}\leq\varphi(t_{i},t_{j})$ (we use the convention that $u_{i,j}:=u_{j,i}$ for $1\leq j<i$).
Let $G=(V,E)$ be a graph as above (i.e. $G=G(\omega,\underline{u})$), where $V$ and $E$ are respectively the vertex set and the edge set. If $\{t,t^{\prime}\}$ is an edge, we write $t\sim t^{\prime}$ and say that $t,t^{\prime}$ are adjacent. A path in $G$ is a string $(t_{1},t_{2},\dots,t_{n})$ such that $t_{k}\in V$ for all $k=1,2,\dots,n$ and $\{t_{k},t_{k+1}\}\in E$ for $k=1,2,\dots,n-1$. Given $A,B,D\subset{\mathbb{R}}^{d}\times{\mathbb{M}}$ we say that there is a path from $A$ to $B$ in $D$ if there is a path $(t_{1},t_{2},\dots,t_{n})$ in $G$ such that $t_{1}\in A$, $t_{n}\in B$ and $t_{k}\in D$ for all $k=1,2,\dots,n$. In this case we write $A\leftrightarrow B$ in DD” or also $A\leftrightarrow B$ in $D$ in GG” when we need to stress that we deal with the graph $G$. If $D={\mathbb{R}}^{d}\times{\mathbb{M}}$, it is omitted from the notation.
@@ 4.2 FKG inequality
To benefit of the FKG inequality we need to switch to a different space. As in [Last17], we define $({\mathbb{R}}^{d}\times{\mathbb{M}})^{[2]}$ as the set $\{A\subset{\mathbb{R}}^{d}\times{\mathbb{M}}\,:\,\sharp A=2\}$. When equipped with the Hausdorff metric this set is a Borel subset of a complete separable metric space [Last17]. We set ${\mathbb{W}}:=({\mathbb{R}}^{d}\times{\mathbb{M}})^{[2]}\times[0,1]$ (trivially, an element of $\mathbb{W}$} is a pair $\big(\{t,t^{\prime}\},u\big)$ with $t\not=t^{\prime}$ in ${\mathbb{R}}^{d}\times{\mathbb{M}}$ and $u\in[0,1]$). We then define $\mathcal{N}({\mathbb{W}})$ as the family of locally finite subsets of $\mathbb{W}$}. Then $\mathcal{N}({\mathbb{W}})$ is a measurable space with $\sigma$–algebra of measurable sets generated by the sets $\{\zeta\in\mathcal{N}({\mathbb{W}})\,:\,\sharp(\zeta\cap B)=n\}$ as $B$ varies among the Borel sets of $\mathbb{W}$} and $n$ varies in $\mathbb{N}$}.
On $\mathbb{W}$}, we define the partial order $\leq$ as follows: $\zeta\leq\zeta^{\prime}$ if for any $(\{t,t^{\prime}\},u)\in\zeta$ it holds $(\{t,t^{\prime}\},u^{\prime})\in\zeta^{\prime}$ for some $u^{\prime}\in[0,u]$. A measurable function $f:\mathcal{N}({\mathbb{W}})\to{\mathbb{R}}$ is called increasing if $f(\zeta)\leq f(\zeta^{\prime})$ whenever $\zeta\leq\zeta^{\prime}$. An event $A$ in $\mathcal{N}({\mathbb{W}})$ is called increasing if $\mathbbm{1}_{A}$ is increasing.
@@
Indeed, the first inequality follows from the FKG inequality for independent uniform random variables with the partial order naturally induced by the standard order on real numbers ($f$ and $g$ are both decreasing functions of the connection variables $u$). The second inequality follows from the FKG inequality for PPPs (cf. [Last18]*Theorem 20.4) as $E[f|\eta]$ and $E[g|\eta]$ are both increasing functions of $\eta$ w.r.t. inclusion (see below), while under $P$ $\eta$ is a PPP with law $P_{\rho,\nu}$. To check the last monotonicity, we observe that, given $\eta,\eta^{\prime}\in\Omega$ with $\eta\subset\eta^{\prime}$, we can realize a coupling between the connection variables associated to $\eta$ and the ones associated to $\eta^{\prime}$ so that for all $x\not=y$ in $\eta\subset\eta^{\prime}$ the connection variable for the pair $\{x,y\}$ is the same for $\eta$ and for $\eta^{\prime}$. Then, the corresponding configurations in $\mathcal{N}({\mathbb{W}})$ are in increasing order w.r.t. $\leq$ (where each configuration is given by the collections of triples $(\{x,y\},u)$, $u$ being the connection variable associated to $\{x,y\}$). The same order is then maintained when applying $f$ or $g$. ∎
@@ 4.3 Irreducibility
For $P_{\lambda,\nu}\otimes{\mathbb{Q}}$–$a.$ $(\omega,\underline{u})$ the following holds: for each $(x,m)\in\omega$ with $m\in A$ the cluster of $(x,m)$ in $G(\omega,\underline{u})$ is finite.
@@
[…]
@@
${\mathbb{P}}(t\leftrightarrow t^{\prime}\leftrightarrow\infty\text{ in }G({\mathbb{X}}\cup\{t,t^{\prime}\}))>0\hskip 18.49988pt\forall(t,t^{\prime})\in({\mathbb{R}}^{d}\times A^{c})^{2}\,.$ (15)
@@
[…]
@@
(we stress that now the expectation is w.r.t. both the PPP $\mathbb{X}$} and the connection variables). By the multivariate Mecke equation (cf. [Last18]*Theorem 4.4), we conclude that for $(\lambda\mathcal{L}\otimes\nu)^{\otimes 2}$ $a.$ $(t,t^{\prime})\in({\mathbb{R}}^{d}\times A^{c})^{2}$ it holds ${\mathbb{E}}[f(t,t^{\prime},{\mathbb{X}}\cup\{t,t^{\prime}\})]=0$, which can be rewritten as ${\mathbb{P}}\left({\mathbb{R}}^{d}\times A\leftrightarrow t\leftrightarrow t^{\prime}\leftrightarrow\infty\text{ in }G({\mathbb{X}}\cup\{t,t^{\prime}\})\right)=0$. Combining this result with (15), we get ${\mathbb{P}}(t\leftrightarrow t^{\prime}\text{ in }G(({\mathbb{X}}\setminus({\mathbb{R}}^{d}\times A))\cup\{t,t^{\prime}\}))>0$ for $(\lambda\mathcal{L}\otimes\nu)^{\otimes 2}$–$a.$ $(t,t^{\prime})\in({\mathbb{R}}^{d}\times A^{c})^{2}$. The last property is equivalent to (3) for $\lambda^{\prime},\nu^{\prime}$.
[…]